# The Riemann Hypothesis is false

### v33

@@
say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
$\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
[6, p.107]. Hence
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{it}=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{T/2<n<T}\Big{)}n^{it}\ll T^{1/2}\log^{3}T.$ (5)
By partial summation, it follows from (5) that
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
@@
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$. For $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$. Thus it follows from (5) by routine partial summation (Riemann-Stieltjes integration) that for such $m$ we have
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (10)
For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it follows again from (5) by partial summation that
@@
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ The author thinks that this approach is not powerful enough to do that. Indeed, Theorem 4.11 of Titchmarsh [6] only works to approximate $\zeta(\sigma+it)$ with a Dirichlet polynomial of length $>Ct/2\pi$ where $C>1$ is a given constant, provided that $\sigma\geq\sigma_{0}>0.$ We therefore cannot apply that result to find a lower bound for $\Theta_{T},$ since $\lim\inf_{T\rightarrow\infty}\Re(1-\rho_{T})=0$.

### v34

@@
say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{n\leq y}n^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}\asymp(t/2\pi)^{1/2-\sigma}$ [6, p.78]. Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
$S(x):=\sum_{n\leq x}n^{it}\ll T^{1/2}\log^{3}T$ (5)
uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}$. Then by partial summation, we obtain from (5) that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
@@
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$. For $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-\frac{1}{\log n}-1}(\log(m)\log(n)-\log^{2}(m)-\log^{2}n)}{n\log^{2}(n)\log^{2}(n/m)}$. Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) by routine partial summation that
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$. For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
@@
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}\sim 0$ with an error term $\ll(\log T)^{-1}$. However, this is not possible (as can be seen from the approximate functional equation for $\zeta$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}\sim 0$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh needs the assumption that $\sigma\geq\sigma_{0}>0$.